Determine the energy required to reverse the direction of a magnetic dipole of 3×10-4 A.m2 in a magmetic field of 0,6 T


1
Expert's answer
2021-05-13T18:02:55-0400

Gives

"M=3\times10^{-4}Am^2"

"B=0.6T"

"\theta=180"

Now

Energy "U=-M.B"

"U= -MBcos\theta"

"cos180\u00b0=-1"

"U=-3\times10^{-4}\times0.6\times cos180°"

"U=1.8\times20^{-4} J"